OpenStudy (anonymous):

2^(x+3)=7 a. -5.19 b.0.94 c. -0.19 d. 2.81

OpenStudy (anonymous):

x=(log7 base 2)-3

OpenStudy (anonymous):

i can get to that point but then i get stuck

OpenStudy (anonymous):

2^3=8 2^n=7 hence n< 3. the answer is negative one. try out more

OpenStudy (zehanz):

You are looking for a number, x+3, that when used as an exponent with base 2 gives 7. So this is the case: \[x+3=\log_{2}7 \].If you wish, you can use the calculator again, to get this log: ln(7)/ln(2), which is about 2.807. Now all you have left to do is subtracting 3 from this number, (because we've found x+3), so the answer must be -0.193. The right answer is c. Do you understand?

OpenStudy (anonymous):

yes thank you